# Don’t Cut Corners: Exact Conditions for Modularity in Biologically Inspired Representations

### v1

@@ Appendix A Related Works
Neuroscience. Theories prompted by the recognition of neural mixed-selectivity in biological networks have argued that these nonlinearly mixed selective codes might exist to enable a linear readout to flexibly decode any required categorisation [50], suggesting a generalisability-flexibility tradeoff between modular and nonlinear mixed encodings [6]. Modelling work has studied task-optimised network models of neural circuits, some of which have recovered mixed encodings [42]. However, other models trained on a wide variety of cognitive tasks have found that networks contain meaningfully modularised components [73, 15, 17]. Dubreuil et al. [16] study recurrent networks trained on similar cognitive tasks and find that, depending on the task structure, neurons in the trained networks can be grouped into populations defined by their connectivity patterns. To this developing body, our work and its predecessor [69] present understanding of when simple biological constraints would encourage encodings of many variables to be modularised into populations of single pure-tuned variables, or when mixed-selectivity would be preferred. Compared to [69] which only works when source variables are statistically independent, our work determines modularity for (almost) any dataset even when source variables are not independent. This is a dramatically more complete understanding of the boundary between mixed and modular codes that allows us to understand behaviour in both biological and artificial neurons that previously eluded explanation.
@@
Independent components analysis and matrix factorization. Our linear autoencoding problem (Theorem 2.1) bears some similarity to linear independent components analysis (ICA) and matrix factorization (MF). These assume data ${\bm{y}}_{i}$ is linearly generated from a set of constrained sources, ${\bm{y}}_{i}={\bm{A}}{\bm{s}}_{i}$, and build algorithms to recover ${\bm{s}}_{i}$ from ${\bm{y}}_{i}$. The literature studies different combinations of (1) assumptions on the sources ${\bm{s}}_{i}$, e.g. independence [5], nonnegativity[49], or boundedness [31]; (2) assumptions on the mixing matrix ${\bm{A}}$, e.g. nonnegative in nonnegative matrix factorisation and (3) algorithms, often deriving identifiability results: conditions that the data must satisfy for a given algorithm to succeed. Some work even derives biological implementations of such factorisation algorithms [47, 10]. In our work we study the conditions under which a particular biologically inspired representation modularises a set of bounded sources. We provide conditions on the sources under which the optimal representation is modular, morally similar to identifiability results (indeed, other identifiability results, like ours, sometimes take the form of sufficient spread conditions [61, 62]). We differ from previous work in the representation we study (which has only been studied before by [69]), the form of our identifiability result (which is both necessary and sufficient, has not been derived before, and assumes only that the sources are bounded), and that these developments allows us to explain both recurrent modularisation and neural data.
@@ B.1.4 Activity Loss and Combination
Let’s analyse the activity loss of these two representations, for the modular representation (eq: 23):
@@ B.1.6 Relaxing Assumptions
Hence, we find that even the loosest inequality possible, eqaution 31, is never satisfied. Further, since the variables are extreme point independent their covariance never achieves the equalities in equation 30, so inequality 31 is not satisfied and the variables should modularise.
@@
Now, assume without loss of generality that for each variable $b_{x}=|\min_{t}x(t)|\leq\max_{t}x(t)$. Then the identical argument to before carries through for the modularisation behaviour around the lower left corner. Around the other corners however, things are more complicated. There might be a corner missing from the lower right quadrant, but to exploit it you would have to build a mixed neuron that is positively tuned to $y$ but negatively tuned to $x$. This would incur a cost, as suddenly the $x$ variable is oriented in its non-optimal direction. Therefore the calculation changes, because the modular solution can always choose to ‘correctly’ orient variables, so doesn’t have to pay this cost. This is all expressed in the same inequalities, eqn. 28:
@@
One quick point is that, for range assymetric variables, the covariance is unbounded, and so it can be large enough to demodularise any pair of variables. Assume for demonstration you have two variables with the same max and min ($b_{\max}$ and $b_{\min}$). The highest covariance you can reach with this pair of bounded, mean-zero, variables is for both variables to spend most of their time either both at the top of their range with some probability $p$, or at the bottom with some probability $1-p$. Because the variables are mean-zero $b_{\max}p=-b_{\min}(1-p)$. Now the covariance is $b_{\min}^{2}\frac{1-p}{p}$, which can grow to any value, and demodularise anything.
@@ B.2 Full Theoretical Treatment
$\operatorname*{minimize}_{{\bm{W}}_{\mathrm{in}},{\bm{b}}_{\mathrm{in}},{\bm{W}}_{\mathrm{out}},{\bm{b}}_{\mathrm{out}}}$ $\quad\left\langle||{\bm{z}}^{[i]}||_{2}^{2}\right\rangle_{i}+\lambda\left(||{\bm{W}}_{\mathrm{out}}||_{F}^{2}+||{\bm{W}}_{\mathrm{in}}||_{F}^{2}\right)$ (35)
@@ B.3 Range-Independent Variables
In the same setting as Theorem 2.1, if $\left|\min_{i}\bar{s}^{[i]}_{j}\right|=\max_{i}\bar{s}^{[i]}_{j}\;\forall j\in[d_{s}]$, i.e. each source is range-symmetric, then the optimal representation modularises if all sources are pairwise extreme-point independent, i.e. if
@@
If sources are range-symmetric and mean-zero then the largest their variance can be is $(\max_{i}s_{j}^{[i]})^{2}$, which occurs when the source takes its maximum and minimum values each with probability $\frac{1}{2}$ [26]. Further, by Cauchy-Schwarz, the covariance between two variables is bounded by their variances, so we get a bound. Further, while two variables are maximally correlated if they are equal, they are maximally anticorrelated if one is minus the other. Hence the covariance is bounded from below by the same quantity with a minus sign. Putting these together:
$-\max_{i}s_{j}^{[i]}\max_{i}s_{k}^{[i]}<\langle s_{j}^{[i]}s_{k}^{[i]}\rangle_{i}<\max_{i}s_{j}^{[i]}\max_{i}s_{k}^{[i]}$ (50)
Create the following vector:
${\bm{m}}=\begin{bmatrix}\max_{i}s_{1}^{[i]}\\ \max_{i}s_{2}^{[i]}\\ \vdots\\ \max_{i}s_{d_{s}}^{[i]}\end{bmatrix}$ (51)
Then, using the bounds on the covariances, we derive that:
$\sum_{j=1}^{d_{s}}\left(w_{j}\min_{i}\bar{s}^{[i]}_{j}\right)^{2}-\sum^{d_{s}}_{j,j^{\prime}\neq j}w_{j}w_{j^{\prime}}\left\langle\bar{s}^{[i]}_{j}\bar{s}^{[i]}_{j^{\prime}}\right\rangle_{i}<({\bm{w}}^{T}{\bm{m}})^{2}$ (52)
Now we just return to the final term in the original inequality. For extreme-point indepedent variables this simplifies considerable:
$(\min_{i}[{\bm{w}}^{\top}{\bm{s}}^{[i]}])^{2}=({\bm{w}}^{T}\min_{i}{\bm{s}}^{[i]})^{2}=({\bm{w}}^{T}{\bm{m}})^{2}$ (53)
Therefore, the inequalities become:
$(\min_{i}[{\bm{w}}^{\top}{\bm{s}}^{[i]}])^{2}=({\bm{w}}^{T}{\bm{m}})^{2}>\sum_{j=1}^{d_{s}}\left(w_{j}\min_{i}\bar{s}^{[i]}_{j}\right)^{2}-\sum^{d_{s}}_{j,j^{\prime}\neq j}w_{j}w_{j^{\prime}}\left\langle\bar{s}^{[i]}_{j}\bar{s}^{[i]}_{j^{\prime}}\right\rangle_{i}$ (54)
Which we can see, using Equation 52, is now true for all ${\bm{w}}$, therefore the representation should be modular, regardless of the correlations. ∎
@@ B.4 Perturbative Approaches to Modularisation of Linearly Mixed Sources
$\operatorname*{minimize}_{{\bm{W}}_{\mathrm{in}},{\bm{b}}_{\mathrm{in}},{\bm{W}}_{\mathrm{out}},{\bm{b}}_{\mathrm{out}}}$ $\quad\left\langle||{\bm{z}}^{[i]}||_{2}^{2}\right\rangle_{i}+\lambda\left(||{\bm{W}}_{\mathrm{in}}||_{F}^{2}+||{\bm{W}}_{\mathrm{out}}||_{F}^{2}\right)$ (55)
@@ B.5.4 Generalisation to Linearly Mixed Sources
B.6 Difficulty of Generalising to Multidimensional Sources
@@
Many variables in the real world, like in our PFC example (Section 5) are not 1-dimensional. Ideally our theory would tell us when such multi-dimensional variables were modularised. Unfortunately, this is not an easy step from our current theory. This is because our proof strategies make extensive use of the optimal modular encoding. For one-dimensional sources this is very easy to find, simply align each source correctly in a neuron, and make the representation positive. Finding similar optimal modular solutions for multidimensional sources is much harder, potentially as hard as simply solving the optimisation problem. We therefore leave it to future work to come up with clever strategies to predict when multidimensional variables should modularise.
Appendix C Positive Linear Recurrent Neural Networks

### v2

@@ Appendix A Related Works
Neuroscience. Theories prompted by the recognition of mixed-selectivity in biological neurons have argued that nonlinearly mixed selective codes might exist to enable a linear readout to flexibly decode any required categorisation [39], suggesting a generalisability-flexibility tradeoff between modular and nonlinear mixed encodings [5]. Modelling work has studied task-optimised network models of neural circuits, some of which have recovered mixed encodings [35]. However, other models trained on a wide variety of cognitive tasks have found that networks contain meaningfully modularised components [59, 12, 14]. Dubreuil et al. [13] study recurrent networks trained on similar cognitive tasks and find that, depending on the task structure, neurons in the trained networks can be grouped into populations defined by their connectivity patterns. To this developing body, our work and its predecessor [56] present understanding of when simple biological constraints would encourage encodings of many variables to be modularised into populations of single pure-tuned variables, or when mixed-selectivity would be preferred. Compared to [56] which only works when source variables are statistically independent, our work determines modularity for (almost) any dataset even when source variables are not independent. This is a dramatically more complete understanding of the boundary between mixed and modular codes that allows us to understand behaviour in both biological and artificial neurons that previously eluded explanation.
@@
Independent components analysis and matrix factorization. Our linear autoencoding problem (Theorem 2.1) bears some similarity to linear independent components analysis (ICA) and matrix factorization (MF). These assume data ${\bm{y}}_{i}$ is linearly generated from a set of constrained sources, ${\bm{y}}_{i}={\bm{A}}{\bm{s}}_{i}$, and build algorithms to recover ${\bm{s}}_{i}$ from ${\bm{y}}_{i}$. The literature studies different combinations of (1) assumptions on the sources ${\bm{s}}_{i}$, e.g. independence [4], nonnegativity[38], or boundedness [26]; (2) assumptions on the mixing matrix ${\bm{A}}$, e.g. nonnegative in nonnegative matrix factorisation and (3) algorithms, often deriving identifiability results: conditions that the data must satisfy for a given algorithm to succeed. Some work even derives biological implementations of such factorisation algorithms [37, 8]. In our work we study the conditions under which a particular biologically inspired representation modularises a set of bounded sources. We provide conditions on the sources under which the optimal representation is modular, morally similar to identifiability results (indeed, other identifiability results, like ours, sometimes take the form of sufficient spread conditions [49, 50]). We differ from previous work in the representation we study (which has only been studied before by [56]), the form of our identifiability result (which is both necessary and sufficient, has not been derived before, and assumes only that the sources are bounded), and these developments allows us to explain both recurrent modularisation and neural data.
Grid Cell Modules One of the many surprising features of grid cells is their modular structure: each grid cell is a member of one module, of which there are a small number in rats. Grid cells from the same module have receptive fields that are translated versions of one another [48]. Previous work has built circuit models of such modules showing how they might path-integrate [9], or has assumed the existence of multiple modules, then shown that they form a good code for space, and that parameter choices such as the module lengthscale ratio can be extracted from optimal coding arguments [34, 54]. However, neither of these directions shows why, of all the ways to code space, a multi-modular structure is best. [11] study a normative problem that explains the emergence of multiple modules grid cells as the best of all possible codes, but their arguments for the emergence of multiple grid modules are relatively intuitive. In this work we are able to formalise parts of their argument.
@@ B.1.4 Activity Loss and Combination
Let’s analyse the activity loss of these two representations, for the modular representation (eq: 24):
@@ B.1.6 Relaxing Assumptions
Hence, we find that even the loosest inequality possible, eqaution 32, is never satisfied. Further, since the variables are extreme point independent their covariance never achieves the equalities in equation 31, so inequality 32 is not satisfied and the variables should modularise.
@@
Now, assume without loss of generality that for each variable $b_{x}=|\min_{t}x(t)|\leq\max_{t}x(t)$. Then the identical argument to before carries through for the modularisation behaviour around the lower left corner. Around the other corners however, things are more complicated. There might be a corner missing from the lower right quadrant, but to exploit it you would have to build a mixed neuron that is positively tuned to $y$ but negatively tuned to $x$. This would incur a cost, as suddenly the $x$ variable is oriented in its non-optimal direction. Therefore the calculation changes, because the modular solution can always choose to ‘correctly’ orient variables, so doesn’t have to pay this cost. This is all expressed in the same inequalities, eqn. 29:
@@ B.2 Full Theoretical Treatment
$\min_{{\bm{W}}_{\mathrm{in}},{\bm{b}}_{\mathrm{in}},{\bm{W}}_{\mathrm{out}},{\bm{b}}_{\mathrm{out}}}$ $\quad\left\langle||{\bm{z}}^{[i]}||_{2}^{2}\right\rangle_{i}+\lambda\left(||{\bm{W}}_{\mathrm{out}}||_{F}^{2}+||{\bm{W}}_{\mathrm{in}}||_{F}^{2}\right)$ (36)
@@ B.3 Range-Independent Variables
In the same setting as Theorem 2.1 the optimal representation modularises if all sources are pairwise extreme-point independent, i.e. if
@@
If the variables are extreme-point independent then we can take the min inside the sum in the definition of the mixed neuron bias:
$z_{n}^{[i]}=\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}+\Delta_{n}=\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}-\min_{i}[\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}]=\sum_{j=1}^{d_{s}}\bigg{(}u_{nj}\bar{s}_{j}^{[i]}-\min_{i}[u_{nj}\bar{s}_{j}^{[i]}]\bigg{)}$ (51)
Now we can compute the activity loss:
$\langle(z_{n}^{[i]})^{2}\rangle_{i}=$ (52)
$\sum_{j=1}^{d_{s}}\langle\bigg{(}u_{nj}\bar{s}_{j}^{[i]}-\min_{i}[u_{nj}\bar{s}_{j}^{[i]}]\bigg{)}^{2}\rangle_{i}+\sum_{j,j^{\prime}=1,j\neq j}^{d_{s}}\langle\bigg{(}u_{nj}\bar{s}_{j}^{[i]}-\min_{i}[u_{nj}\bar{s}_{j}^{[i]}]\bigg{)}\bigg{(}u_{nj}\bar{s}_{j^{\prime}}^{[i]}-\min_{i}[u_{nj}\bar{s}_{j^{\prime}}^{[i]}]\bigg{)}\rangle_{i}$
The cross terms at the end are strictly positive, and each of the squared terms on the left is the same or larger than its corresponding modular neuron, since either $u_{nj}$ is positive and it is the same, or $u_{nj}$ is negative and it is greater than or equal to the equivalent modular neuron. Therefore the mixed energy cost is larger than the modular energy cost for extreme-point independent variables. ∎
@@ B.5.4 Generalisation to Linearly Mixed Sources
Appendix C Extension to all Activity Norms
We now extend our theoretical results to the same setting as the main theorem, Theorem B.1, but with other activity norms rather than L2. As shown in theorem Theorem B.2, the weight loss is minimised by orthogonalising the coding of different variables, so we just focus on the activity loss and will recombine later. We will show that, for all norms, modular solutions are optimal if the variables are range independent.
Consider a set of modular neurons:
${\bm{z}}_{n}^{[i]}=\begin{bmatrix}|u_{n1}|(\bar{s}_{1}^{[i]}-\min_{j}\bar{s}_{1}^{[j]})\\ \vdots\\ |u_{nd_{s}}|(\bar{s}_{d_{s}}^{[i]}-\min_{j}\bar{s}_{d_{s}}^{[i]})\end{bmatrix}\geq{\bf 0}$ (91)
And let’s see if we could usefully mix these encodings together for any set of mixing coefficients ${\bm{u}}_{n}\in\mathbb{R}^{d_{s}}$:
$z_{n}^{[i]}=\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}+\Delta_{n}=\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}-\min_{i}[\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}]\geq 0$ (92)
If the variables are range-independent then we can move the min inside the sum:
$\Delta_{n}=-\min_{i}[\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}]=-\sum_{j=1}^{d_{s}}\min_{i}u_{nj}\bar{s}_{j}^{[i]}$ (93)
Then the Lp norm is:
$\langle|z_{n}^{[i]}|^{p}\rangle_{i}=\langle|\bigg{(}\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}+\Delta_{n}\bigg{)}|^{p}\rangle_{i}=\langle|\bigg{(}\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}-\min_{i}u_{nj}\bar{s}_{j}^{[i]}\bigg{)}|^{p}\rangle_{i}$ (94)
Now using the convexity of the Lp norm:
$\langle|\bigg{(}\sum_{j=1}^{d_{s}}u_{nj}\bar{s}_{j}^{[i]}-\min_{i}u_{nj}\bar{s}_{j}^{[i]}\bigg{)}|^{p}\rangle_{i}\geq\sum_{j=1}^{d_{s}}\langle|u_{nj}\bar{s}_{j}^{[i]}-\min_{i}u_{nj}\bar{s}_{j}^{[i]}|^{p}\rangle_{i}$ (95)
with equality only if p is equal to 1, or if at most one of the entries of ${\bm{u}}$ are 0.
Finally, since we orient our variables such that $|\min_{i}\bar{s}^{[i]}_{j}|<\max_{i}\bar{s}^{[i]}$, without loss of generality,:
$\langle|u_{nj}\bar{s}_{j}^{[i]}-\min_{i}u_{nj}\bar{s}_{j}^{[i]}|^{p}\rangle_{i}\geq\langle|u_{nj}|^{p}\langle|\bar{s}_{j}^{[i]}-\min_{i}\bar{s}_{j}^{[i]}|^{p}\rangle_{i}$ (96)
So we get our final result that the mixed activity loss is greater than or equal to the modular activity loss if the variables are range independent:
$\langle|z_{n}^{[i]}|^{p}\rangle_{i}\geq\sum_{j=1}^{d_{s}}|u_{nj}|^{p}\langle|\bar{s}_{j}^{[i]}-\min_{i}\bar{s}_{j}^{[i]}|^{p}\rangle_{i}=\langle|{\bm{z}}_{n}^{[i]}|^{p}\rangle_{i}$ (97)
So, as long as $p>1$, the activity loss is lowered by modularising the representation of range-independent variables. Further, since the weight loss is also minimised, the optimal solution will be modular.
Two caveats: first, in the particular case of the L1 norm all that we can guarantee is that the activity loss of a modular solution is equal to that of a mixed solution. This means that, even in the case of range-independent variables, there might be orthogonal but mixed solutions that are equally good as the modular solution. For $p>1$ this caveat is not needed.
Second, this is a set of sufficient conditions, range-independent variables are optimally encoded modularly. There will be other nearly range-independent variables that are also optimally encoded modularly, just as we found for the L2 case.
Appendix D Extension to Multidimensional Variables
@@
Many variables in the real world, such as angles on a circle, are not 1-dimensional. Ideally our theory would tell us when such multi-dimensional sources were modularised. Unfortunately, precise necessary and sufficient conditions are not an easy step from our current theory because our proof strategies make extensive use of the optimal modular encoding. For one-dimensional sources this is very easy to find, simply align each source correctly in a neuron, and make the representation positive. Finding similar optimal modular solutions for multidimensional sources is much harder, potentially as hard as simply solving the optimisation problem.
However, we can make progress on a simpler problem: showing that range-independence is a sufficient condition for a set of multi-dimensional sources to modularise. Consider two multidimensional sources ${\bm{x}}^{[i]}\in\mathbb{R}^{d_{x}}$ and ${\bm{y}}^{[i]}\in\mathbb{R}^{d_{y}}$, and an encoding in which they are mixed in single neurons:
$z_{n}^{[i]}=\sum_{j=1}^{d_{x}}u_{nj}x_{j}^{[i]}+\sum_{j=1}^{d_{y}}v_{nj}y_{j}^{[i]}+\Delta_{n}$ (98)
Whatever choice of coefficient vectors, ${\bm{u}}$ and ${\bm{v}}$, or activity norm (as long as $p>1$), we can consider breaking this mixed coding into two modular neurons. Define new variables $a^{[i]}=\sum_{j=1}^{d_{x}}u_{nj}x_{j}^{[i]}$ and $b^{[i]}=\sum_{j=1}^{d_{y}}v_{nj}y_{j}^{[i]}$. If the two multidimensional sources are range-independent, then so too are $a$ and $b$. Define the modular representation:
${\bm{z}}_{n}^{[i]}=\begin{bmatrix}a^{[i]}-\min_{j}a^{[j]}\\ b^{[i]}-\min_{j}b^{[j]}\end{bmatrix}$ (99)
All of our previous arguments apply to argue that this modular encoding will be preferred if the multidimensional variables are range-independent.
We just have to show that the weight loss is also reduced by modularising multidimensional sources, a small extension of Theorem B.2, which we shall now show. Write a modular code as:
${\bm{z}}^{[i]}={\bm{U}}{\bm{x}}^{[i]}-\min_{j}[{\bm{U}}{\bm{x}}^{[j]}]+{\bm{V}}{\bm{y}}^{[i]}-\min_{j}[{\bm{V}}{\bm{y}}^{[j]}]$ (100)
With the encoding of ${\bm{x}}$ and ${\bm{y}}$ spanning two different spaces: ${\bm{U}}^{T}{\bm{V}}={\bf 0}$.
Then define the minimal L2 norm readout weight matrix using the pseudoinverse:
${\bm{W}}_{\text{out}}=\begin{bmatrix}{\bm{U}}^{\dagger}\\ {\bm{V}}^{\dagger}\end{bmatrix}$ (101)
Such that ${\bm{W}}_{\text{out}}{\bm{z}}^{[i]}=\begin{bmatrix}{\bm{x}}^{[i]}\\ {\bm{y}}^{[i]}\end{bmatrix}$ up to some constant offset that can be removed by the bias term in the readout. Now consider mixing the representation of the two multidimensional sources using an orthogonal matrix that preserves the encoding size of each variable, ${\bm{O}}$:
${\bm{z}}^{[i]}={\bm{U}}{\bm{x}}^{[i]}+{\bm{O}}{\bm{V}}{\bm{y}}^{[i]}\qquad{\bm{W}}_{\text{out}}={\bm{W}}_{\text{out}}=\begin{bmatrix}{\bm{U}}^{\dagger}+{\bm{A}}\\ {\bm{V}}^{\dagger}{\bm{O}}^{T}+{\bm{B}}\end{bmatrix}$ (102)
Then in the subspace spanned by the columns of ${\bm{U}}$ ${\bm{W}}_{\text{out}}$ must be the same as ${\bm{U}}^{\dagger}$. However, in order to offset the newly aligned matrix ${\bm{O}}\tilde{{\bm{V}}}$ there must be an additional component orthogonal, ${\bm{A}}$:
${\bm{W}}_{out}{\bm{U}}={\bm{U}}^{\dagger}{\bm{U}}+{\bm{A}}{\bm{U}}=\mathbb{I}\qquad{\bm{A}}{\bm{U}}={\bf 0}$ (103)
${\bm{W}}_{out}{\bm{O}}{\bm{V}}={\bm{U}}^{\dagger}{\bm{O}}{\bm{V}}+{\bm{A}}{\bm{O}}{\bm{V}}={\bf 0}\qquad{\bm{A}}{\bm{O}}{\bm{V}}=-{\bm{U}}^{\dagger}{\bm{O}}{\bm{V}}$
So ${\bm{A}}$ and ${\bm{U}}^{\dagger}$ live in orthogonal subspaces, and ${\bm{A}}\neq 0$. The same applies to ${\bm{V}}$ and ${\bm{B}}$, therefore:
$||{\bm{W}}_{out}||_{F}^{2}=||{\bm{U}}^{\dagger}||_{F}^{2}+||{\bm{V}}^{\dagger}||_{F}^{2}+||{\bm{A}}||_{F}^{2}+||{\bm{B}}||_{F}^{2}\geq||{\bm{U}}^{\dagger}||_{F}^{2}+||{\bm{V}}^{\dagger}||_{F}^{2}$ (104)
And hence orthogonalising the subspaces reduces the loss in this case as well.
Combining these two, the optimal representation of range-independent multi-dimensional variables for activity losses using an Lp norm for $p>1$ is modular.
Appendix E Positive Linear Recurrent Neural Networks